A battery lasts, average 4 years with a standard deviation of 1.4 years.the battery life is normally distributed

1
Expert's answer
2021-12-13T08:28:25-0500

Let "X="the lifetime of a battery: "X\sim N(\mu, \sigma^2)"

Given "\mu=4, \sigma=1.4"

a)


"P(X<3)=P(Z<\dfrac{3-\mu}{\sigma})=P(Z<\dfrac{3-4}{1.4})"

"\approx P(Z<-0.7143)\approx0.2375"

b)


"P(1.8<X<5)=P(X<5)-P(X\leq1.8)"

"=P(Z<\dfrac{5-4}{1.4})-P(Z\leq\dfrac{1.8-4}{1.4})"

"\approx P(Z<0.7143)-P(Z\leq -1.5714)"

"\approx0.76248-0.05804\approx0.7044"


c)


"P(X<x)=P(Z<\dfrac{x-4}{1.4})=0.25"

"\dfrac{x-4}{1.4}\approx-0.6745"

"x\approx4-1.4(0.6745)"

"x\approx3.0557"